AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 552     Entry time: Fri May 10 21:23:35 2024
Author: TD, CB, JM 
Subject: 20Ne7+ test beam 

22.00

Carsten Brandau stopping e-target experiment Unattended beam overnight. No operators watching.

10^5 (nominal) 20Ne7+ at 380keV/u beam available

Carsten reports beam is unstable, often disappearing, and in bunched mode.

4x MSL type TTT14 DSSSDs not moving, locked in out position

Starting CARME MIDAS DAQ ASIC settings 2024Apr24-14.14.03

slow comparator 0x64 E

G&G Ortec 448 Pulser OFF

DSSSD bias & leakage currents - attachment 1

FEE64 temperatures OK - attachment 2

All system wide checks OK, WR timestamps OK - attachment 3

ADC data item stats - attachment 4

Rates dominated by two hot channels aida15 per FEE64 Stat spectra - attachment 5

per FEE64 1.8.L spectra - attachments 6-7 no obvious beam energy peak, continuum of energies from below beam energy to threshold

per FEE64 1.8.W spectra - attachments 8-9

 

22.55 Start data file APR24/R5 Merger, TapeServer etc - attachments 10-12

270k data items/s

 

23:29 Activated CARMESort script. See attachments 13-15.

No obvious beam or signal of any kind. Temperatures OK.

 

00:13 DAQ continues OK. System-wide check OK. Temps OK. Still no clear sign of beam.

 

01:03 Daq continues OK. System wide checks OK, Temps OK.

Only low energies observed in the energy spectrum. Looks a bit like the low energy noise observed in our previous runs with no/very few counts above what would be beam energy (~7.6 MeV)

Note the xy plots are not all orientated correctly (strips are correct in the sort code but root axis need to be flipped - I can;t remember how atm). The part of the plots with counts are those closest to the beam. 

02:00 DAQ continues OK. System wide checks OK. Temps OK.

Spectra still looks the same. I have noticed the right hand side rate is higher than the left. EyEx cut spectrum also looks like much more counts outside of the cut.

03:40

DAQ continues. All checks and temps OK.

I can now see a small peak at 5 MeV in the bottom left dssd. This dssd also has the least number of counts in the low energy continuum.

I have logged back into grafana on x2go to get the number ions in the ring. Attachment 17

04:23 (05:23 EU)

DAQ stopped, merger stopped, tape server stopped. Bias OFF. FEEs turned OFF. Screen shots taken of the monitoring code plots - not yet uploaded.

Windows and servers remain up in case we want to do some more tests in the near future.

Attachment 18 shows beam intensity for total run time from Grafana

Attachment 1: Screenshot_from_2024-05-10_22-31-48.png  39 kB  Uploaded Fri May 10 22:42:11 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-31-48.png
Attachment 2: Screenshot_from_2024-05-10_22-32-52.png  146 kB  Uploaded Fri May 10 22:42:25 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-32-52.png
Attachment 3: Screenshot_from_2024-05-10_22-33-45.png  292 kB  Uploaded Fri May 10 22:42:47 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-33-45.png
Attachment 4: Screenshot_from_2024-05-10_22-34-59.png  136 kB  Uploaded Fri May 10 22:43:06 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-34-59.png
Attachment 5: Screenshot_from_2024-05-10_22-37-12.png  219 kB  Uploaded Fri May 10 22:43:22 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-37-12.png
Attachment 6: Screenshot_from_2024-05-10_22-38-32.png  160 kB  Uploaded Fri May 10 22:43:41 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-38-32.png
Attachment 7: Screenshot_from_2024-05-10_22-39-48.png  165 kB  Uploaded Fri May 10 22:43:54 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-39-48.png
Attachment 8: Screenshot_from_2024-05-10_22-40-38.png  269 kB  Uploaded Fri May 10 22:44:09 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-40-38.png
Attachment 9: Screenshot_from_2024-05-10_22-41-24.png  264 kB  Uploaded Fri May 10 22:44:23 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-41-24.png
Attachment 10: Screenshot_from_2024-05-10_22-56-53.png  530 kB  Uploaded Fri May 10 22:57:21 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-56-53.png
Attachment 11: Screenshot_from_2024-05-10_22-56-20.png  125 kB  Uploaded Fri May 10 22:57:37 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-56-20.png
Attachment 12: Screenshot_from_2024-05-10_22-56-35.png  122 kB  Uploaded Fri May 10 22:57:54 2024  | Hide | Hide all
Screenshot_from_2024-05-10_22-56-35.png
Attachment 13: Screenshot_from_2024-05-10_23-28-37.png  424 kB  Uploaded Fri May 10 23:33:19 2024  | Hide | Hide all
Screenshot_from_2024-05-10_23-28-37.png
Attachment 14: Screenshot_from_2024-05-10_23-28-45.png  288 kB  Uploaded Fri May 10 23:33:30 2024  | Hide | Hide all
Screenshot_from_2024-05-10_23-28-45.png
Attachment 15: Screenshot_from_2024-05-10_23-28-51.png  451 kB  Uploaded Fri May 10 23:33:38 2024  | Hide | Hide all
Screenshot_from_2024-05-10_23-28-51.png
Attachment 16: Screenshot_from_2024-05-11_00-12-43.png  280 kB  Uploaded Sat May 11 00:14:34 2024  | Hide | Hide all
Screenshot_from_2024-05-11_00-12-43.png
Attachment 17: 20240511_025056.jpg  3.686 MB  Uploaded Sat May 11 04:26:33 2024  | Hide | Hide all
20240511_025056.jpg
Attachment 18: 20240511_042236.jpg  4.426 MB  Uploaded Sun May 12 15:35:15 2024  | Hide | Hide all
20240511_042236.jpg
ELOG V3.1.4-unknown